OpenStudy (anonymous):

Can anyone help me out? got some algebra 2..

OpenStudy (anonymous):

What is the solution set to the following set? x+y=4 x^2 + y^2=16

rishavraj (rishavraj):

see its x + y = 4 on squaring both sides \[\(x \+\ y)\^\2\ \=\ \1\6\\] so x^2 + y^2 + 2xy = 16

rishavraj (rishavraj):

and u r given tht x^2 + y^2 = 16 i.e the second equation so u get 2xy = 0

rishavraj (rishavraj):

@Cman456

OpenStudy (anonymous):

(4,0) (0,-4)?

rishavraj (rishavraj):

and wht about (4,0) hey it can't be negative,......its given x + y =4

rishavraj (rishavraj):

sorry i meant (0,4)

OpenStudy (anonymous):

so( 4,0) (0,4)

rishavraj (rishavraj):

yeah.....

OpenStudy (anonymous):

Another one?

OpenStudy (anonymous):

Y=(x-3)(x+2)(x-2) what are the zeroes of the function?

OpenStudy (anonymous):

-3,2.-2?

rishavraj (rishavraj):

its supposed to be 3, -2, 2

OpenStudy (anonymous):

um why

OpenStudy (anonymous):

nope nvm got it

OpenStudy (anonymous):

another one please haha

OpenStudy (anonymous):

classify 4x^5 + 2x^4 -5x^3+12 by number of terms

OpenStudy (anonymous):

polynomial of 4 terms?

OpenStudy (anonymous):

or trinomial?

OpenStudy (bee_see):

4 terms.

OpenStudy (bee_see):